OpenStudy (anonymous):

Solve: x^2 +2y^2 =33 3x +2y = -11 The answer is suppose to be: ( -5,2) and (-1,-4) but idk how they got that! :(

OpenStudy (anonymous):

helllppp! ;o

OpenStudy (whpalmer4):

Do you recognize what those formulas are?

OpenStudy (anonymous):

yeah they are for Conic solutions. and they are inequalities right?

OpenStudy (whpalmer4):

The first one is an ellipse, the second one is a line.

OpenStudy (anonymous):

Oh yeah that i know.

OpenStudy (whpalmer4):

the solutions are the points where the line crosses the ellipse.

OpenStudy (anonymous):

Yeha, i know how to do it by graphing but I need to know how to solve it by using equations etc

OpenStudy (whpalmer4):

Right, I'm just making sure you understand the setup. Here's how you can solve it exactly: Take the second equation and solve it for one of the variables. Say you solve it for \(x\): now you replace \(x\) in the first equation with whatever you got when you solved the second equation for \(x\), which will give you an equation in terms of \(y\). Solve that for \(y\) — it will be a quadratic and give you two solutions. Plug those two values of \(y\) into either original equation to find the corresponding values of \(x\).

OpenStudy (anonymous):

Okay i did everytihng up till hte quadratic part but its not factorable . so i dont know how to get the 2 answers.

OpenStudy (whpalmer4):

what do you get for a quadratic?

OpenStudy (anonymous):

i got till this point and then get confused. x^2 +2 (.5(-11-3x))^2 = 33

OpenStudy (anonymous):

because i get like decimal points after this and no whole numbers

OpenStudy (whpalmer4):

You solved for \(y\) and got \(y = \frac{1}{2}(-11-3x)\) Generally better to avoid using decimals when doing algebra... \[x^2 + 2(\frac{1}{2}(-11-3x))^2 = 33\]Now we just carefully expand that \[x^2+2*(\frac{1}{2})^2(-11-3x)^2 =33\]\[x^2+\frac{1}{2}((-11)^2+33x+33x+9x^2) = 33\] Try it from there...

OpenStudy (whpalmer4):

Might be convenient to multiply everything by 2 to get rid of that fraction...

OpenStudy (whpalmer4):

and then divide everything by 11 once you've done that

OpenStudy (whpalmer4):

I think you'll find that you can factor the result.

OpenStudy (anonymous):

i got 10x^2+66x+88

OpenStudy (anonymous):

is that right?

OpenStudy (whpalmer4):

No, that's not correct. for starters, it isn't an equation.

OpenStudy (anonymous):

aww what did i do wrong?

OpenStudy (whpalmer4):

I'm not sure, you didn't show me your work...

OpenStudy (anonymous):

x^2 +.5 (121+33x+33x+9x^2) = 33 10x^2+66x+88=0

OpenStudy (whpalmer4):

what did I say about using decimals while doing algebra? By doing that instead of writing a fraction, you completely missed it when you expanded :-( \[x^2+\frac{1}{2}((-11)^2+33x+33x+9x^2) = 33\]Multiply by 2: \[2x^2 + 2*\frac{1}{2}((-11)^2+33x+33x+9x^2) = 2*33\]\[2x^2+(121+66x+9x^2) = 66\]

OpenStudy (whpalmer4):

collect like terms, divide out any common factors, then factor for the solution...

OpenStudy (anonymous):

AHHH ur the best! thank you!

OpenStudy (anonymous):

i got it!

OpenStudy (whpalmer4):

Great!